OpenStudy (anonymous):

A bag contains 3 green marbles and 8 white marbles. Suppose a marble is randomly selected. What are the odds in favor of picking a white marble?

OpenStudy (anonymous):

number of white marbles / total number of marbles

OpenStudy (anonymous):

is it okay?

OpenStudy (anonymous):

8/11

hartnn (hartnn):

that is not odds in favor, that is probability

OpenStudy (anonymous):

8/11 is not right answer because i got that problem wrong

hartnn (hartnn):

that is because 8/11 is probability, not what u wanted

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so what is it?

hartnn (hartnn):

let probability be p = 8/11 here odds in favor = p/(1-p)

hartnn (hartnn):

which comes out to be 8/3

OpenStudy (anonymous):

thats what i thought

OpenStudy (anonymous):

@hartnn good one bro if you are a male..or sis if you are a girl

OpenStudy (anonymous):

I missed out "odds"!!! pellet man

OpenStudy (anonymous):

The probability of winning a particular game at the Game Place is 1/4. Find the odds against winning the game. can you guys help me with this.

hartnn (hartnn):

odds against = (1-p)/p p= 1/4 here

OpenStudy (anonymous):

okay,

OpenStudy (anonymous):

its not 3 to 4 because i gave that answer and it said it was wrong

hartnn (hartnn):

because u calculated it incorrectly, it comes out to be 3/1 = 3

hartnn (hartnn):

(3/4)/(1/4)=3

OpenStudy (anonymous):

ohhhhh

OpenStudy (anonymous):

thannkks

hartnn (hartnn):

welcomes ^_^

OpenStudy (anonymous):

Remember that concept...it is useful in many problems..

OpenStudy (anonymous):

21 is 35% of what number? lol anyhelp with this too since your so smart (:

hartnn (hartnn):

let 21 be 35% of x x% = x/100 so 35x/100 = 21 solve this

OpenStudy (anonymous):

(35/100)x = 21 Solve

OpenStudy (anonymous):

lol..its already said

hartnn (hartnn):

type faster bro

OpenStudy (anonymous):

wait do i X the 35 and 100 or divide?

OpenStudy (anonymous):

yoohooo.. I will do it the next time

OpenStudy (anonymous):

because i think i divide

hartnn (hartnn):

35x/100 = 21 multiply 100 on both sides 35x= 2100

hartnn (hartnn):

now divide 35 on both sides

OpenStudy (anonymous):

60

hartnn (hartnn):

yup

OpenStudy (anonymous):

yep..that's right @dianarock336

hartnn (hartnn):

lol, late again :P

OpenStudy (anonymous):

okay thannks lol you guys figghht alot.

hartnn (hartnn):

O.o

OpenStudy (anonymous):

no no I just did it for fun..coz he typed faster than me!!!